OpenStudy (anonymous):

...a junior one day admission to an amusement park is $23.this price is increased by 75% for adults.find the mark up and the adult price.

OpenStudy (anonymous):

easiest way to increase something by 75% is just multiply it by 1.75

OpenStudy (anonymous):

Multiply it by 0.75 then add it to the 23.

OpenStudy (anonymous):

0.75 * 23 = 17.25 + 23 = ?

OpenStudy (anonymous):

40.25?

OpenStudy (anonymous):

it is asking for two prices,

OpenStudy (anonymous):

Or...like i said 3 hours ago...23 x 1.75...gives you the same answer.